# SGLDBench: A Benchmark Suite for Stress-Guided Lightweight 3D Designs

### v1

@@ 5 Conclusion and Outlook
SGLDBench is a benchmark suite for stiff lightweight design simulation and analysis, with special emphasis on high-resolution 3D designs. It supports the computation of six different types of designs, using methods spanning the wide spectrum from purely stiffness-based optimization to geometry-aware infill generation. SGLDBench enables users to create different types of designs at different resolutions and with different material consumptions, and to assess their mechanical and geometric properties. Direct volume rendering with ambient occlusion is used to improve the visual perception of high-resolution 3D designs. SGLDBench visualizes the deviations between stress directions in the initial solid object and a generated design to reveal relationships between stiffness, geometric properties, and stress replication.
SGLDBench can be used to compute individual designs with specific boundary conditions. New design strategies can be integrated into the publically available code basis. Over that, SGLDBench can be used to create new types of designs by combining existing design strategies. For instance, the material field of an optimal design can be computed via topology optimization and downloaded to extract the design boundary as an isosurface in this field. This surface can then be uploaded to SGLDBench to compute an infill that is restricted to the surface’s interior. An example is shown in the supplemental material.
By using SGLDBench’s functionality, a number of interesting properties of existing design strategies have been revealed, guiding toward new research directions in lightweight design: Why is the mechanical performance of conforming lattice structures below expectation in 3D domains? What is the role of adaptive porosity in achieving high stiffness? What is the interplay between truss thickness, cell size, and stiffness in lattice infills? How does an optimal smoothing operator for a tensor field look so that the stiffness deviation is minimized, e.g., so that forces are still transmitted along the most efficient load paths? How does a Voronoi infill perform if further constraints on the edges regarding stress reproduction are enforced? SGLDBench supports researchers in finding answers to these questions and developing improved design strategies regarding different objectives.
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG) under grant number WE 2754/10-1. Several open-source projects are used when developing SGLDBench, including the volumetric Michell structure [11], conforming lattice structures [50], the tetrahedral mesh generator TetGen [56], and the voxelization functions [36], the MMA implementations [40, 57]. Hereby, We sincerely acknowledge the generosity of the authors for making their code publicly available.

### v2

@@ 5 Conclusion and Outlook
SGLDBench is a benchmark suite designed for the simulation and analysis of stiff, lightweight designs, with a special emphasis on high-resolution 3D models. It supports the computation of six distinct design types, employing methods that range from purely stiffness-based optimization to geometry-aware infill generation.
SGLDBench enables users to create designs at varying resolutions and material consumption levels while assessing their mechanical and geometric properties. Various visualization options provide additional insights into governing stress states and a design’s geometric structure. The resistance of a design to new force situations can be assessed via the color coding of designs with scalar stress measures. Additionally, SGLDBench visualizes deviations between stress directions in the initial solid object and the generated design, providing insights into relationships between stiffness, geometric properties, and stress replication.
The suite allows users to compute individual designs with specific boundary conditions and offers flexibility for integrating new design strategies into its publicly available code base. Moreover, SGLDBench supports the creation of novel design types by combining existing strategies. For instance, the material field of an optimized design can be computed via TO and downloaded to extract its boundary as an isosurface. This surface can then be uploaded to SGLDBench to compute an infill restricted to the interior of the surface, as demonstrated in the supplemental material.
By leveraging SGLDBench’s capabilities, several intriguing properties of existing design strategies have been uncovered, paving the way for new research directions in lightweight design. For example, why does the mechanical performance of conforming lattice structures fall below expectations in 3D domains, what role does adaptive porosity play in achieving high stiffness, what is the interplay between truss thickness, cell size, and stiffness in lattice infills, how can a tensor field be optimally smoothed to minimize stiffness deviation while maintaining efficient load transmission paths, or how would a Voronoi infill perform if additional constraints on edge stress reproduction were applied? SGLDBench provides researchers with tools to investigate these questions and develop improved design strategies tailored to diverse objectives.
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG) under grant number WE 2754/10-1. Several open-source projects are used when developing SGLDBench, including the volumetric Michell trusses [35], conforming lattice structure [6], tetrahedral mesh generators TetGen [59] and fTetWild [60], voxelization functions [37], and MMA implementations [41, 61]. Hereby, we sincerely acknowledge the generosity of the authors in making their code publicly available.